### v1

# Analyzing Classroom Interaction Data Using Prompt Engineering and Network Analysis

Abstract
Classroom interactions play a vital role in developing critical thinking, collaborative problem-solving abilities, and enhanced learning outcomes. While analyzing these interactions is crucial for improving educational practices, the examination of classroom dialogues presents significant challenges due to the complexity and high-dimensionality of conversational data. This study presents an integrated framework that combines prompt engineering with network analysis to investigate classroom interactions comprehensively. Our approach automates utterance classification through prompt engineering, enabling efficient and scalable dialogue analysis without requiring pre-labeled datasets. The classified interactions are subsequently transformed into network representations, facilitating the analysis of classroom dynamics as structured social networks. To uncover complex interaction patterns and how underlying interaction structures relate to student learning, we utilize network mediation analysis. In this approach, latent interaction structures, derived from the additive and multiplicative effects network (AMEN) model that places students within a latent social space, act as mediators. In particular, we investigate how the gender gap in mathematics performance may be mediated by students’ classroom interaction structures.
Keywords: Classroom Interaction, Prompt Engineering, Network Analysis, Additive and Multiplicative Effect Network Model, Network Mediation Analysis
1 Introduction
1.1 Background: Classroom Dialogue Analysis
Classroom interactions play a crucial role in the learning process and substantially influence students’ motivation, comprehension, and academic achievement. Through effective student-teacher and student-student interactions, learners develop essential skills in critical thinking, collaborative work, and problem-solving. Analysis of these student interactions reveals underlying behavioral patterns that either facilitate or impede effective learning.
The analysis of classroom dialogues provides a powerful lens for examining student interactions, offering a direct understanding of engagement patterns and learning processes. Sedova et al. (2019) demonstrates a significant correlation between participation in substantive dialogue and improved academic performance. Similarly, Mercer (2010) establishes that student conversations foster collaborative problem-solving abilities and critical thinking skills through collective meaning-making. These empirical findings underscore the dual nature of classroom dialogues: they simultaneously manifest student interactions and serve as the primary medium through which these interactions can be systematically investigated.
Conventional Approaches
While classroom conversations offer valuable insights into learning dynamics, their analysis presents considerable analytical challenges due to their inherent complexity and high dimensionality. These conversations are often characterized by dynamic shifts in topic, speaker turns, and the interplay of various communicative functions, making manual analysis particularly difficult. These also encompass speaker characteristics (e.g., their speaking style, experience), emotional tone (e.g., enthusiasm, uncertainty, affect), and context (e.g., type of task, classroom environment). To address such complexity, researchers often simplify the analysis by categorizing the intent of individual messages or utterances using predetermined coding schemes. For example, researchers may classify utterances according to function, such as whether they are questions, answers, elaborations, or agreements (Webb et al., 2014; Franke et al., 2015). This strategy entails classifying utterances based on semantic, syntactic, and pragmatic features, and using these classifications to gain insights into specific aspects of the interaction.
[…]
These limitations highlight the need for automated analytical strategies to thoroughly examine classroom dynamics. Although several deep learning methods have been proposed for automatic dialogue classification (Liu et al., 2017; Raheja and Tetreault, 2019; Song et al., 2021), they depend on large labeled datasets that are difficult to acquire and may not generalize well across different research objectives. For example, Raheja and Tetreault (2019) proposed a context-aware self-attention model for general dialogue act classification, but it requires extensive manual annotation for adaptation to specific tasks. On the other hand, Song et al. (2021) proposed an automated model for classroom dialogue classification by combining a convolution neural network (CNN) and a bidirectional-long short-term memory (Bi-LSTM), but it also requires large labeled datasets and may have difficulty generalizing to diverse research objectives. These challenges motivate the development of alternative approaches that minimize reliance on pre-labeled datasets and offer greater flexibility.
Prompt Engineering
[…]
Strategic prompt engineering enables researchers to classify dialogue interaction types even without traditional training data for fine-tuning, provided the prompts incorporate sufficient contextual information. Nevertheless, while prompt engineering facilitates classification of individual utterances, it cannot fully capture the complex dynamics of classroom interactions. The classification of discrete utterances reveals speaker identities and discussion content but fails to illuminate critical patterns such as interaction frequency, structural relationships, and the relationships between students.
Network Analysis
To address these gaps, network analysis offers a powerful approach for analyzing the structure of student interactions (Grunspan et al., 2014; Bokhove, 2018). By representing the classroom as a network, with students as nodes and their interactions as edges, researchers can analyze complex patterns of communication frequency and type within broader classroom dynamics. For example, centrality metrics can reveal key insights about student influence and prominence within the network (Williams et al., 2019). To uncover deeper patterns of local interactions and latent social dynamics, latent space modeling techniques can be used (Hoff et al., 2002; Hoff, 2005; Sosa and Buitrago, 2021). The projection of students into a latent social space facilitates visualization and analysis of underlying classroom social structures, revealing both student clusters with similar interaction patterns and the influence of these patterns on classroom dynamics and learning outcomes. Network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022) can be used to further investigate how classroom network structures influence student learning outcomes. By using latent vectors from latent space modeling, this approach quantifies the mechanisms through which classroom interactions shape achievement, offering insights into how social structures mediate educational outcomes.
1.2 Purpose and Structure of the Paper
In this paper, we present an integrated framework for analyzing classroom interactions by combining the strengths of prompt engineering and network analysis. Our methodology begins with automated dialogue classification through prompt engineering, addressing the limitations of traditional manual labeling approaches. We employ carefully constructed prompts to guide the generative model in classifying interactions according to predefined criteria, establishing the foundation for network-based analysis of classroom dynamics. We then transform these classified interactions into network representations and analyze them using a network mediation model. Specifically, we employ the additive and multiplicative effects network (AMEN) model (Hoff, 2021) model to extract latent social structures and utilize them as mediators to investigate how classroom networks mediate the relationship between gender and mathematics performance. Through examination of how these latent social structures influence student outcomes, with particular attention to gender differences in mathematics performance, we illuminate the mechanisms by which network structures mediate the impact of classroom dynamics on student achievement. This comprehensive methodological integration can provide educational researchers with a powerful analytical tool for investigating the complex relationships between social dynamics and academic performance.
The remainder of this paper is structured to present our methodology and findings systematically. Section 2 examines our prompt engineering approach, detailing the automated utterance classification process and the creation of labeled interaction data. Section 3 introduces our network analysis framework, beginning with the conceptual foundations of the network mediation model, progressing through the application of the AMEN model for latent position estimation, and concluding our key analytical findings. Section 4 concludes by exploring the broader implications of our research and identifying promising directions for future studies that build upon this integrated methodological framework.
2 Prompt Engineering for Utterance Classification
2.1 Data: Classroom Interaction Dataset
We analyze classroom interaction data from Webb et al. (2023). The dataset includes observations of 20 third-grade students (9 male, 11 female) from a classroom located in a low-income urban community in Southern California. Researchers collected data over a five-month period, during which time they recorded six one-hour math lessons using strategically placed cameras to capture comprehensive classroom interactions between teachers and students, as well as interactions among students. The recordings were carefully transcribed to include contextual information, including speaker identities, timestamps, and job specifications. Initially, students’ mathematics performance was assessed with the California standards test (CST) (EdData, 2013). Due to the statewide transition in the assessment framework, however, the final performance was assessed by the author-developed multiplication and division test (Webb et al., 2023). The pre-test, based on CST, yielded a mean score of 369.35 (SD = 54.31). The post-test, specifically designed to assess problem-solving strategies and accuracy, resulted in a mean score of 15.08 out of 24 (SD = 6.49).
[…]
A significant challenge in analyzing dialogue data arises from the context-dependent nature of utterances. The meaning of conversations often goes beyond individual statements, necessitating a careful examination of the surrounding dialogue. The following brief dialogue example illustrates how seemingly simple responses can convey subtle meanings when considered within their conversational context:
Student A: I think it’s 39 because each one had 39.
Student B: Did you put it one by one?
Student A: Yeah.
Student C: Mine got 35.
[…]
2.2 Prompt Engineering Methodology
The complexity and high-dimensional nature of classroom interactions necessitate a robust method for classifying and analyzing dialogue intent. As discussed in the Introduction Section, existing analytical approaches require manual labeling, which is resource-intensive, while deep learning methods require extensive labeled datasets. Our methodology addresses these issues through prompt engineering, by leveraging pre-trained models to classify dialogue intents accurately without labeled training data or computational fine-tuning. This approach maximizes the capabilities of pre-trained models while maintaining efficiency in resource utilization and processing time.
In-context learning (ICL) is a specialized application of prompt engineering in which pre-trained large language models (LLMs) learn exclusively from information provided within the prompt, eliminating the need for parameter updates. This approach parallels human learning processes through inference while minimizing computational requirements for task adaptation (Dong et al., 2022). Few-shot prompting, one of the ICL strategies, enhances model comprehension by incorporating targeted task examples within the prompt. The effectiveness of few-shot prompts in ICL has been explained through probabilistic frameworks such as hidden Markov models (Baum and Petrie, 1966), where carefully selected examples guide model reasoning by establishing sequential dependencies that reduce prediction uncertainty (Xie et al., 2021). Min et al. (2022) further demonstrates that optimal performance depends not merely on increasing example quantity but on developing prompt formats that precisely align with task-specific requirements, including both structural elements and linguistic formulation.
[…]
Figure 1: System and User Prompts for Utterance Classification. Includes guidelines, category definitions, and example dialogues for analyzing and classifying the contribution of utterances in collaborative problem-solving processes.
Figure 1 illustrates the user and system prompts implemented in this study. The prompts define the utterance classification task, outline categorization rules, and provide illustrative dialogues to clarify distinctions between categories. The user prompt provides clear guidelines for classifying utterances into three main categories: ‘explain own ideas’, ‘engage others’ ideas’, or ‘uncorrelated’. For the ‘engage others’ ideas’ category, it further defines three levels of engagement (low, medium, high) and offers specific criteria for making classification decisions, including special cases and exceptions. The system prompt supports these guidelines by stressing the importance of accurate categorization, emphasizing subtle differences between labels, and ensuring consistency across classifications. To enhance output reliability and reduce variability in the model’s responses, we adjusted the temperature parameter to nearly zero, optimizing consistency in classification.
2.3 Utterance Classification Results
We conducted utterance classification using prompts carefully constructed, as illustrated in Figure 1. These prompts were designed incorporating comprehensive contextual information, including dialogue tasks, categorization rules, and classification guidelines. To optimize classification accuracy, we included five few-shot examples in the prompts, each structured to specify speaker identity, utterance content, classification label, engagement level, and relevant references. Complete prompt specifications, including the few-shot examples, are detailed in the Section C of the Supplementary Material. Instead of individual utterances, dialogue chunks segmented by timestamps from the original dataset were included as the inputs for classification within our prompts. This approach allows the prompts to capture conversational context effectively and inform the classification decisions. The classification process was executed on an NVIDIA RTX A6000 GPU with an Intel(R) Xeon(R) Gold 6426Y CPU in a Linux environment, requiring approximately 1.5 hours for completion.
[…]
  Utterance Type Engagement Level Number of Utterances
Explain Own Idea (EXP) 1,335
Engage Others Idea (EOI) Total 1,251
High Engagement 107
Medium Engagement 553
Low Engagement 591
Uncorrelated 1,544
Total Utterances 4,130
Table 1: Classification Results by Utterance Type and Engagement Level
[…]
  Name texts Label Engagement Level
  Delilah Okay now. First I put 2 days to represent the days. And then I put 70 units in each group (gestures across the two groups on her notebook). explain own idea -
Keila First I… uncorrelated -
Delilah And I came up with 39 because each one had 39. explain own idea -
Keila Did you put it one by one? engage others idea medium
Delilah Yeah. explain own idea -
Keila Mine got 35. engage others idea low
Delilah So I got 1, 2, 3, 4, 5, 6, 7, 8, blah blah blah. (counting the marks in one group) explain own idea -
Keila Did you put 1, 2, 3, 4, 5, 6? (gestures marking in alternating groups as she counts, demonstrating action in solving) engage others idea medium
Delilah Yeah! And 10 divided by 2 is 3. I meant 39. explain own idea -
Table 2: Examples: Identifying labels (explain vs. engage) and engagement levels.
  Name texts Label Engagement Level
  William Okay so first what I did with this is I did 70 and 2. So I had two days. 1 day, 2 days, and I counted all the way up to 70 by 2, and I counted. And so now I got 32. explain own idea -
Julian I disagree with you cause I got 1, 2, 3, (…), 14 (counts his 14 rods) engage others idea high
William Wait hold on. You don’t add 7 plus 7. If you add 7 plus 7 that would be 14. But are you trying to add? Or times? engage others idea high
Table 3: High-Level Engagement in Problem-Solving
Second, Table 3 illustrates the system’s ability to capture complex interaction sequences. The example begins with a student’s explanation of their approach (classified as EXP), followed by a peer’s substantive disagreement (classified as high-level EOI). The subsequent probing question, which builds upon this disagreement, maintains the high-level EOI classification. This sequence demonstrates the capacity of the system to identify sophisticated interactions where students progressively build upon and critically engage with each other’s mathematical reasoning.
  Name texts Label Engagement Level
  Natalie Okay, let me go. uncorrelated -
Kimberly I missed a number. uncorrelated -
Natalie I’m looking at the hundreds chart. (refers to yellow sheet) And then I went to 75. explain own idea -
Kimberly Why did you go to 75? engage others idea medium
Natalie Because that’s how much shells… explain own idea -
Kimberly (talks over N) No, it was 70. engage others idea low
Natalie …that was how much shells she found. explain own idea -
Kimberly It was 70. engage others idea low
Natalie On the second strategy. explain own idea -
Kimberly Oh! uncorrelated -
Natalie And then, I was thinking how could I do it. Then, I was thinking about money, I don’t know why. Then, I remembered a quarter, so then I know what that is. (pause) And then, I started like because I remember the quarters, so I went 20, 25, no… explain own idea -
Kimberly (talking to another group) You should keep on sharing strategies. That’s what we are doing. uncorrelated -
Samantha We are already done. uncorrelated -
Kimberly We are at the strategies, but she and me realized that we strategies on the 200 chart. uncorrelated -
Table 4: Differentiating Relevant and Off-Topic Remarks.
[…]
2.4 Validation of Utterance Classification
For the validation of our prompt engineering approach, we focused on assessing the alignment between intended classifications and actual outcomes. Due to the absence of pre-labeled data in our classroom dialogue dataset, we utilized the DailyDialog dataset (Li et al., 2017) as our external validation source. This dataset provides a structured framework for dialogue act classification, categorizing utterances into four distinct types: inform, questions, directives, and commissive. To evaluate our methodology, we developed dialogue act classification (DAC) specific prompts for this dataset, incorporating five dialogue chunks examples to create few-shot examples prompts. We compared the performance of these few-shot prompts against zero-shot approaches, enabling us to quantify the impact of including dialogue examples in the classification process.
To establish the effectiveness and generalizability of our prompt engineering approach, we conducted comprehensive comparisons with established deep learning models. Context-Aware Self-Attention model (Raheja and Tetreault, 2019, CASA) employs a self-attention mechanism integrated with hierarchical deep neural networks to recognize dialogue acts within conversational contexts. The Deep Co-Interactive Relation Network (Qin et al., 2020, DCR-Net) employs a hierarchical encoder and relation layer to integrate contextual and interaction information for dialogue act classification. A co-interactive graph attention network model (Qin et al., 2021, Co-GAT) jointly performs DAC and sentiment classification, leveraging both contextual and interaction information through cross-utterance and cross-task connections. Additionally, Zhao et al. (2023) presented classification results using prompt engineering with the GPT-3.5-turbo model, and we compared our results against these presented results. We evaluated classification performance using three complementary metrics: macro-F1, weighted-F1 scores, and overall accuracy. Macro-F1 computes the F1 score for each class individually and then averages them, providing a balanced measure of performance across all classes, regardless of their frequencies. This is particularly useful in imbalanced datasets, where certain classes may dominate. On the other hand, Weighted-F1 calculates the F1 score for each class and takes into account the number of true instances for each class, giving more importance to classes with more instances. These metrics help assess the model’s performance from different perspectives, considering both class balance and overall accuracy.
  Model Acc Macro-F1 Weighted-F1
  CASA - 0.78 -
DCR-Net + Co-Attention - 0.79 -
Co-GAT - 0.79 -
ChatGPT 1-shot + P.E 0.67 - 0.65
ChatGPT 3-shot + P.E 0.73 - 0.71
Llama-3.1-8B-Instruct Zero-shot 0.69 0.57 0.68
Llama-3.1-8B-Instruct Few-shot 0.75 0.66 0.74
Table 5: Comparison of Prompt Engineering results with other baselines on the DAC task. Experimental results for CASA, DCR-Net + Co-Attention, and Co-GAT are partly cited from Qin et al. (2021) and Zhao et al. (2023). P.E. stands for prompt engineering where the 1-shot and 3-shot settings use 1 and 3 examples in the prompt, respectively.
Table 5 presents our comparative analysis of prompt engineering approaches and state-of-the-art deep learning models. Results demonstrate that the few-shot prompt engineering method, incorporating five dialogue chunks as examples, achieves substantially improved performance compared to the zero-shot approach. While some comparative studies did not report all evaluation metrics, we have calculated a comprehensive set of metrics for our prompt engineering results to facilitate thorough comparison across different models. This analytical approach ensures equitable performance assessment across varying methodologies.
The comparative analysis reveals that prompt engineering performance, while effective, may not fully match that of deep learning models due to fundamental methodological distinctions. Deep learning approaches benefit from extensive optimization using pre-constructed labeled datasets and generate specialized embeddings tailored to specific tasks. In contrast, prompt engineering operates solely through label descriptions without prior training data, relying on the inherent language understanding capabilities of the model. This distinction becomes particularly significant in contexts where classification criteria exhibit variability due to multiple annotators or crowdsourcing processes. Nevertheless, these differences suggest that prompt engineering may be especially valuable in generating consistent labels, as it helps maintain classification criteria across datasets.
3 Network Analysis of Classroom Interaction
3.1 Network Construction from Classified Utterances
After classifying utterances using prompt engineering, we created a directed network representation to systematically analyze student interaction patterns. The representation of student interaction patterns as networks enables both visualization and systematic analysis of classroom dynamics.
We formalized student interactions using an adjacency matrix, where the matrix elements indicate the frequency of specific interaction types between two students during the observation period. We then visualized the constructed network structure by using nodes to represent individual students and directed weighted edges between nodes to represent the nature and frequency of the interactions between the two students. For instance, when a student offers multiple explanations to a peer, a directed edge is formed from the explaining student (source) to the listening student (target), with the edge weight reflecting the total number of explanatory interactions. Engagement interactions similarly generate directed edges from the engaging student to the idea originator, with edge weights reflecting interaction frequency. To capture the varying intensity of student engagement more precisely, we implemented a differential weighting scheme for these interactions. While explanatory interactions maintained equal weights, we applied multiplication factors to different levels of engagement: medium-level engagements received double weight, and high-level engagements received triple weight relative to low-level engagements. This weighted approach ensures that more substantive forms of engagement exert proportionally greater influence in the network structure, better reflecting their educational significance.
While this network construction process allows us to represent valuable aspects of classroom interactions, analyzing classroom conversation data in a network also presents unique analytical challenges. Unlike conventional social network datasets, our classroom interactions are often characterized by limited and sparse connections, as conversations frequently occur within small groups during specific question discussions. To overcome the sparsity of individual lesson networks and facilitate a more comprehensive analysis of overall interaction patterns, we aggregated data across all six lessons to construct a single composite network.
3.1.1 Network Graphs
Figure 2 illustrates these network transformations, where each node represents a student and edges represent their interactions. Node colors indicate students’ standardized mathematics performance, utilizing a spectrum from blue (lower scores) to red (higher scores). Edge thickness corresponds to interaction frequency between student pairs.
(a) Explain Own Idea Network, Pre-test
(b) Explain Own Idea Network, Post-test
(c) Engage Others’ Idea Network, Pre-test
(d) Engage Others’ Idea Network, Post-test
Figure 2: Explain Own Idea and Engage Others’ Idea Networks. Edge weights in each network represent the frequency of corresponding utterances, and node colors represent math test scores (red = higher, blue = lower). The left-hand figures show pre-test results, and the right-hand figures show post-test results. The networks are constructed using data aggregated from six lessons.
The network visualization in Figure 2 reveals that interaction quality, rather than quantity, plays a more significant role in influencing learning outcomes. For instance, Julian and William, who initially demonstrated lower mathematics performance in the pre-test, established strong connections primarily with high-achieving peers within the network. Their subsequent improvement in post-test scores suggests that interactions with high-performing students may create multiplicative learning benefits. Conversely, Mya’s case provides an instructive contrast: despite maintaining frequent peer interactions, her academic performance showed minimal improvement. This pattern reinforces the observation that the qualitative aspects of student interactions — specifically, the academic performance levels of interaction partners — may exert greater influence on learning outcomes than the mere frequency of exchanges.
Network Centrality
We then examined network centrality, using metrics designed for directed weighted networks, accounting for both directional flow and interaction intensity. These include PageRank (Page et al., 1999), hub-authority scores (Kleinberg, 1999), in-degree and out-degree, betweenness (Freeman, 1977), closeness (Bavelas, 1948), and eigenvalue centrality (Bonacich, 1972).
Each centrality measure captures distinct aspects of student engagement in the network. PageRank evaluates student influence by considering both the number and importance of incoming connections, assigning higher scores to students who receive interactions from influential peers. Hub scores identify students who initiate connections with multiple peers, while authority scores highlight students who serve as primary information sources. In-degree and out-degree centrality measure the weighted volume of received and initiated interactions respectively, identifying key participants in collaborative discourse. Eigenvalue centrality assesses each student’s influence by considering the influence levels of their connected peers.
Betweenness and closeness centrality measures require a different approach to weight calculation, as they interpret edge weights as distances rather than strengths. We employed the inverse of utterance frequencies as weights, ensuring that frequent interactions correspond to shorter distances. Consequently, betweenness centrality identifies students who serve as key intermediaries in information flow, while closeness centrality quantifies each student’s overall proximity to peers in the interaction network. This inverse weighting scheme ensures that stronger connections (more frequent interactions) contribute to higher centrality scores.
  Pagerank Indegree Outdegree Betweenness Closeness Eigen Hub Authority
  Angelina 0.1033 57 61 32 0.3695 0.2245 0.0378 0.0654
Armando 0.0414 46 46 34 0.4829 0.2242 0.1974 0.0198
Daniel 0.0286 48 37 27 0.4728 0.1811 0.0216 0.6029
Diego 0.0451 75 47 115 0.5616 0.5980 0.0147 0.7355
Esther 0.0479 92 105 76 0.5176 0.8422 0.0556 0.6152
Gio 0.0515 56 139 145 0.5454 0.2639 1.0000 0.0171
Isabel 0.0330 42 40 29 0.4775 0.2533 0.2659 0.0570
Jessica 0.0522 71 26 57 0.4916 0.7276 0.0133 0.5414
Julian 0.0530 87 51 52 0.5046 0.3090 0.0385 0.9111
Keila 0.0229 35 22 0 0.3728 0.1329 0.0080 0.0877
Kevin 0.0484 89 55 71 0.4571 0.2615 0.0045 1.0000
Kimberly 0.0481 51 93 51 0.4613 0.2644 0.1562 0.0420
Marco 0.0518 49 73 18 0.3276 0.1015 0.0067 0.2131
Maria 0.0607 41 53 61 0.4259 0.2476 0.1405 0.0570
Melissa 0.0184 25 46 0 0.3872 0.2847 0.1741 0.0217
Mya 0.0665 92 135 142 0.5677 1.0000 0.6806 0.0618
Natalie 0.0343 60 37 4 0.4005 0.4354 0.0421 0.2712
Ruben 0.0481 59 74 43 0.4086 0.1663 0.5060 0.0051
Samantha 0.0965 77 38 5 0.3629 0.2925 0.0247 0.0741
William 0.0485 62 36 7 0.3242 0.1290 0.1091 0.0133
Table 6: Centrality measures for the EXP (Explain Own Idea) network.
  Pagerank Indegree Outdegree Betweenness Closeness Eigen Hub Authority
  Angelina 0.0307 48 48 36 0.2968 0.0045 0.0006 0.0030
Armando 0.0431 96 50 9 0.6094 0.4061 0.0796 0.0262
Daniel 0.0358 36 33 0 0.4462 0.1607 0.0215 0.0174
Diego 0.0556 96 128 135 0.6766 0.5410 0.0292 0.4539
Esther 0.0964 192 75 110 0.6310 1.0000 0.0162 1.0000
Gio 0.0744 134 75 125 0.5837 0.3982 0.0477 0.0443
Isabel 0.0377 73 38 17 0.5011 0.2453 0.0999 0.0430
Jessica 0.0278 59 80 90 0.5569 0.1406 0.0098 0.1212
Julian 0.0600 71 152 57 0.5309 0.2287 0.0521 0.0290
Keila 0.0228 33 40 0 0.4228 0.1101 0.0055 0.0623
Kevin 0.0615 75 86 38 0.4894 0.2484 0.0398 0.0203
Kimberly 0.0537 73 97 63 0.5287 0.2807 0.2482 0.0173
Marco 0.0707 104 54 10 0.4116 0.2906 0.0092 0.0144
Maria 0.0275 51 77 68 0.5223 0.0788 0.0449 0.0060
Melissa 0.0481 59 58 11 0.5133 0.3470 0.3616 0.0274
Mya 0.0646 119 175 167 0.6586 0.6136 1.0000 0.0229
Natalie 0.0329 47 67 0 0.4311 0.1986 0.0909 0.1434
Ruben 0.0629 82 76 29 0.4609 0.2486 0.0209 0.0447
Samantha 0.0276 37 47 0 0.2647 0.0019 0.0044 0.0002
William 0.0662 71 100 20 0.4313 0.2460 0.0153 0.0144
Table 7: Centrality measures for the EOI (Engage Others’ Idea) network.
Tables 6 and 7 illustrate how different centrality measures reveal distinct patterns of student engagement in the interaction networks. For example, in the explanation network, Mya demonstrates high PageRank centrality, indicating her frequent interactions with influential peers. Esther’s elevated in-degree and out-degree scores reflect her active participation as both recipient and initiator of mathematical explanations. Notably, Gio’s high betweenness centrality suggests a crucial role in facilitating information flow between different student groups, effectively bridging distinct classroom communities.
Analysis of Table 7 reveals distinct interaction patterns within the Engagement network, with several students demonstrating notable centrality profiles. Esther’s consistently high scores across centrality metrics indicate her comprehensive engagement in mathematical discourse, both influencing and responding to peers’ ideas. Gio maintains his bridging role while actively participating in peer discussions, while Mya demonstrates strong bilateral engagement through both contributing to and building upon others’ mathematical reasoning. Additionally, Armando, Diego, and Julian’s elevated eigenvector centrality and hub scores suggest their significant influence within localized peer groups, indicating the presence of distinct collaborative clusters within the classroom.
Our analysis shows that high network centrality does not necessarily correlate with improved mathematical performance. Mya’s case illustrates this complexity: despite her prominent position in both interaction networks, her mathematical achievement did not improve, possibly due to her frequent interactions with lower-performing peers. This suggests that the relationship between network centrality and academic performance is not straightforward and the specific composition of local interactions - such as which students interact with whom - appears to play a more significant role. The influence of classroom interactions on learning outcomes appears to depend more on the composition of immediate peer groups than on students’ overall network prominence. This observation underscores the importance of examining both the quality and context of student interactions within local network structures, rather than relying exclusively on global centrality metrics.
3.2 Network Mediation Model
Following the descriptive analysis of classroom interaction networks through visualization and centrality metrics, we now proceed to a more rigorous examination of how these network structures influence student outcomes. While network centrality measures offer valuable insights into the roles and prominence of individual students within the interaction network, they do not directly reveal the mechanisms through which these interaction patterns impact learning. To understand these mechanisms, we employ network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022). Network mediation analysis provides a framework for addressing this challenge by treating the network structure as a mediating variable. A key challenge in implementing network mediation analysis is representing the complex network structure in a way that can be used as a mediator. To address this, we adopt the approach of latent space model (Hoff et al., 2002; Hoff, 2005; Sewell and Chen, 2016, LSM)). LSM provide a framework to represent each student’s position within the network in a lower-dimensional latent space, capturing their underlying social dynamics wherein inter-student distances reflect the probability and intensity of their interactions.
Figure 3: Multiple mediator model for network mediation analysis
Figure 3 presents a diagram of a multiple mediator model for network mediation analysis. In this model, $M_{1}$ through $M_{k}$ represent the latent positions of network nodes embedded in a latent space. The paths $a_{11}$ through $a_{1k}$ denote the effects of $X$ on the mediators, while the paths $b_{21}$ through $b_{2k}$ capture the effects of the mediators on $Y$. The direct effect of $X$ on $Y$ is represented by path $b_{1}$. By linking these latent positions to independent and dependent variables through regression frameworks, we can decompose total effects into direct and indirect components mediated by the underlying network structure.
Network as a Mediator for Gender Gap in Mathematics Performance
Our study examines how classroom interaction networks mediate the influence of gender on mathematics achievement. Prior research has consistently investigated gender differences in mathematics performance, revealing complex patterns and influences (Lindberg et al., 2010; Else-Quest et al., 2010; Ghasemi and Burley, 2019), while Lindberg et al. (2010) demonstrate that overall gender gaps in mathematics performance are relatively small. Ghasemi and Burley (2019) identifies significant cultural and societal influences on mathematics-related preferences and self-confidence across genders. Else-Quest et al. (2010) emphasizes the role of societal context in gender disparities, noting smaller performance differences in countries with greater gender equality. These findings collectively suggest that gender differences in mathematics achievement stem from social rather than purely biological factors, underscoring the importance of analyzing classroom social networks and their mediating effects on academic performance.
Based on this perspective, we hypothesize that the gender gap in mathematics performance ($X$ on $Y$) can be differentiated based on how individual students interact with others during mathematics discussion classes. Specifically, we apply network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022), using students’ interaction structures as a mediator ($M$) for gender ($X$) - math performance ($Y$) relationships. By incorporating the estimated latent positions ($z_{i}$, $w_{j}$) as mediators, our goal is to investigate the indirect effects of gender on mathematics performance through the network structure.
Pre-test Scores as a Confounder
In the current context, pre-test scores could serve as a useful confounder for the intended mediation analysis. However, the pre-test (California Standardized test; CST) and post-test (a test developed by Webb et al. (2023)) do not measure the same domains, while the assessment methods and criteria are also different between the two tests. Thus, the two test scores may not be entirely comparable (Liu et al., 2017).
Figure 4: Scatter plot showing the relationship between pre-test (CST) and post-test scores.
Further, there is an additional technical challenge: the pre-test scores do not seem to have a linear relationship with the post-test scores. As shown in Figure 4, the relationship is not clearly linear around the Basic (300-350) and Proficient (350-400) ranges of the pre-test. This means that higher scores in the pre-test may not predict higher post-test scores consistently, suggesting pre-test scores may not be a reliable predictor of the post-test scores.
To account for students’ prior mathematical proficiency in our analysis, we constructed a binary proxy variable based on California Standards Test (CST) performance. Students who achieved scores of 350 or above — the state-designated threshold for proficiency [^2] https://www.starsamplequestions.org/about.html — were categorized as high-performing. This dichotomization aligns with state educational standards and provides a meaningful distinction in student achievement levels. We incorporate this proxy variable as a confounder in our mediation analysis to more precisely evaluate how network structures mediate gender-based differences in mathematics performance.
Figure 5: Network mediation model path diagram. This model analyzes the effect of gender on math achievement mediated by network interactions. Latent positions of sender and receiver networks are used as mediators.
Network Mediation Model
We additionally consider predictor-mediator interaction effects to account for potential heterogeneity in mediator effects across varying predictor levels (VanderWeele and Vansteelandt, 2009; Valeri and VanderWeele, 2013). In this case, the indirect effects, mediated through the latent network structure, vary with gender ($x$). By modeling this interaction, we account for the possibility that the indirect effect of gender on math performance is mediated differently depending on the students’ positions within the network. The resulting network mediation model can be expressed as:
$\begin{split}E(M_{k}\mid x,c)&=a_{0k}+a_{1k}\,x+a_{2k}\,c,\\ E(Y\mid x,\mathbf{m},c)&=b_{0}+b_{1}\,x+\sum_{k}b_{2k}\,m_{k}+\sum_{k}b_{3k}\,x\,m_{k}+b_{4}\,c,\end{split}$ (1)
where $x$ is the binary indicator for gender, $m_{k}$ represents the latent positions of students, and $c$ is the confounder indicating proficiency based on the CST score threshold. Specifically, the latent variables $m_{k}$ represent the coordinates of the latent positions for the network nodes: $m_{1}$ and $m_{2}$ correspond to sender position coordinates (x and y axes, respectively), while $m_{3}$ and $m_{4}$ represent receiver position coordinates (x and y axes, respectively). The model coefficients capture distinct aspects of the gender-performance relationship: $a_{1k}$ quantifies gender’s influence on network positioning, $b_{2k}$ measures how network positions affect mathematics performance, and $b_{3k}$ captures the predictor-mediator interaction between gender and network position, revealing how network-mediated effects vary by gender. Additionally, $a_{2k}$ and $b_{4}$ account for the influence of prior mathematical proficiency on network positions and post-test performance, respectively. $a_{0k}$ and $b_{0}$ are intercept terms for the network mediation model.
This model enables the decomposition of the total effect (TE) into two distinct components: the natural direct effect (NDE) and the natural indirect effect (NIE). The NDE represents the component of the effect of gender ($X$) on math performance ($Y$) that is independent of the mediators, while the NIE captures the effect transmitted through the mediating network structure. These components constitute the total effect (TE), providing a comprehensive measure of the impact of gender on math performance. The natural direct effect, the natural indirect effect, and the total effect, are given by:
$\begin{split}NDE&=\left(b_{1}+\sum_{k}b_{3k}\left(a_{0k}+a_{1k}\,x^{*}+a_{2k}\,c\right)\right)(x-x^{*}),\\ NIE&=\left(\sum_{k}\left(b_{2k}\,a_{1k}+b_{3k}\,a_{1k}\,x\right)\right)(x-x^{*}),\\ TE&=NDE+NIE.\end{split}$ (2)
These effects can be calculated from the regression coefficients estimated with Equation (1).
Our analysis examines how distinct types of classroom interactions mediate gender-based differences in mathematics performance by investigating two separate networks: the EXP network and the EOI network. For both networks, we incorporate students’ prior mathematics achievement as a confounder to account for pre-existing academic differences. Through the application of our network mediation model to these distinct interaction networks, we aim to investigate whether the mediating effects of classroom network structures vary systematically based on interaction type.
3.3 AMEN Model for Latent Position Estimation
To obtain the latent positions that serve as mediators, we employ the AMEN model (Hoff, 2021). The AMEN model is an extension of latent space model (LSM; Hoff et al., 2002), designed to capture more sophisticated network patterns, particularly in directed networks like classroom interactions. This extension is particularly relevant for our analysis of classroom interactions, which are inherently directed. The AMEN model explicitly models directed relationships by providing distinct latent positions for each student as both a sender and a receiver. The AMEN model can be expressed as:
$g(\mathbb{E}(y_{ij}))=\alpha_{i}+\beta_{j}+z_{i}^{T}w_{j},$ (3)
where $y_{ij}$ is the value representing the connection between the $i$th node and the $j$th node (e.g., the number of interactions, a binary indicator of whether an interaction occurred), $\alpha_{i}$ and $\beta_{j}$ represent sender and receiver random effects, respectively, and $z_{i}$ and $w_{j}$ are low-dimensional latent feature vectors for nodes $i$ as a sender and $j$ as a receiver respectively, in a $d$-dimensional latent space, where $d=2,3$ is typically considered for visualization purposes. $g(\cdot)$ is a link function. The additive components, $\alpha_{i}$ and $\beta_{j}$, account for individual node-specific effects, while the multiplicative term, $z_{i}^{T}w_{j}$, captures interactions between latent vectors. This approach allows the model to represent complex third-order network dependencies, such as transitivity, clustering, and degree heterogeneity.
AMEN with a Negative Binomial Regression
A unique feature of classroom interaction data is its count-based nature, reflecting the frequency of student interactions. By incorporating these interaction frequencies as edge weights in the network structure, we intend to capture both the presence and intensity of student-to-student connections.
For count data, a Poisson link function is typically considered. However, a Poisson model often suffers from overdispersion, where the observed variance exceeds the theoretical mean. The sender and receiver random effects can account for overdispersion to some degree by allowing the conditional mean to vary across participants (Gelman, 2007). However, severe overdispersion may not be fully addressable with the sender and receiver random effects.
In our analysis of classroom interaction data, we identified a severe overdispersion issue, as detailed in Section D of the Supplementary Material. To effectively address overdispersion, we propose using a negative binomial regression model, which introduces a dispersion parameter to explicitly account for the excess variability and allows the conditional variance to vary independently of the mean. By combining the negative binomial specification with sender and receiver random effects from the AMEM model, we can effectively capture both individual-level heterogeneity and the overdispersed nature of classroom interaction counts.
The proposed AMEN model with a negative binomial regression structure for count data can be written as:
$\log(\mu_{ij})=\alpha_{i}+\beta_{j}+z_{i}^{T}w_{j},$ (4)
where $\mu_{ij}$ is the mean of the negative binomial distribution where $y_{ij}\sim\mbox{NB}(\mu_{ij},r)$ and $r$ of the negative binomial distribution controls the variance of the distribution, with higher values indicating greater dispersion relative to the mean. $\alpha_{i}$ and $\beta_{j}$ are sender and receiver random effects, respectively, and $z_{i}$ and $w_{j}$ are sender and receiver latent feature vectors, respectively.
3.4 Estimation Methods
We take a two-stage estimation approach for data analysis: (1) AMEN model estimation; and (2) network mediation model estimation.
AMEN Model Estimation
In the first stage, we estimate the proposed AMEN model to embed the observed student interaction network into a latent space. The prior distributions of the AMEN model parameters are specified as follows:
$\begin{split}\alpha_{i}&\sim N(0,\sigma_{\alpha}^{2}),\quad\sigma_{\alpha}^{2}>0,\quad i=1,\dots,N\\ \beta_{j}&\sim N(0,\sigma_{\beta}^{2}),\quad\sigma_{\beta}^{2}>0,\quad j=1,\dots,N\\ z_{i}&\sim N_{d}(0,\sigma_{z}^{2}I_{d}),\quad\sigma_{z}^{2}>0,\quad i=1,\dots,N\\ w_{j}&\sim N_{d}(0,\sigma_{w}^{2}I_{d}),\quad\sigma_{w}^{2}>0,\quad j=1,\dots,N\\ \log r&\sim N(\mu_{r},\sigma_{r}^{2}),\quad\mu_{r}\in\mathbb{R},\quad\sigma_{r}^{2}>0\end{split}$ (5)
where $\sigma_{\alpha}^{2}$, $\sigma_{\beta}^{2}$, $\sigma_{z}^{2}$, $\sigma_{w}^{2}$, and $\sigma_{r}^{2}$ represent fixed prior variance terms for each corresponding parameter, and $I_{d}$ denotes the $d\times d$ identity matrix. In this study, we set $d=2$ for interpretability and visualization purposes.
We draw the posterior samples of the AMEN model parameters and latent positions using a Metropolis-Hastings-within-Gibbs sampler. Additional details on the prior means, standard deviations, and the jumping rule for each parameter, as well as the starting points and overall convergence checks, are provided in the Section A of the Supplementary Material.
Note that the inner product between latent vectors remains invariant under transformations such as rotation or reflection. To address this, we perform Procrustes matching (Gower, 1975) to align the latent vectors to a common reference position. This alignment ensures that the latent vectors are consistently positioned, facilitating a clearer interpretation of their relationships and enhancing the validity of subsequent analyses.
Network Mediation Model Estimation
In the second stage, we estimate a network mediation model. Specifically, the latent positions from the estimated AMEN model serve as mediators in the regression framework outlined in Equation (2). The network mediation model is estimated through Bayesian regression, with prior distributions specified as:
[…]
where $\Theta\in\{a_{01},\dots,a_{2k},b_{0},b_{1},b_{21}\dots,b_{3k},b_{4}\}$ represents all regression coefficients in Equation (1), $\epsilon_{i}$ represents the error term in the regression model, and $\sigma^{2}$ is the variance of error term. The network mediation model is estimated using the MCMCpack package (Martin et al., 2011) in R. The estimated regression coefficients are then used to calculate the natural direct effect (NDE) and natural indirect effect (NIE), as defined in Equation (2).
3.5 Network Analysis Results
3.5.1 AMEN Model Results
(a) EXP (Explain Own Idea) Network
(b) EOI (Engage Others’ Idea) Network
Figure 6: Comparison of Latent Positions for EXP and EOI Networks. Each node is represented as a point in a two-dimensional space, where the direction from the origin indicates similarity between nodes, and the distance from the origin indicates influence within the network.
Latent Position Structure
Figures 6(a) and 6(b) illustrate the latent positions of students within both the EXP and EOI networks through two-dimensional visualizations. In these representations, each student’s position is mapped as a point in two-dimensional space. The angle between student vectors indicates their interaction similarity - smaller angles represent greater similarity in interaction patterns. Additionally, vector magnitude represents each student’s relative influence within the network structure.
Figure 2(a) reveals important bridging roles within the network structure, particularly through the positions of Gio and Isabel. Gio functions as a critical connector between two distinct student subgroups: one comprising Diego, Daniel, and Armando, and another consisting of Julian, William, and Marco. Similarly, Isabel facilitates connections between the subgroup of Melissa, Natalie, and Esther and the subgroup of Julian, William, and Marco. The latent vector visualization in Figure 6(a) captures these bridging relationships effectively, positioning Gio centrally between the connected subgroups while placing Isabel in closer proximity to the subgroup of Melissa, Natalie, and Esther.
In Figure 2(c), Isabel serves a bridging role in connecting two distinct student clusters: the Keila and Kimberly group and the Marco and Julian group. In Figure 6(b), Isabel appears closer to the Keila, Kimberly, Melissa, Natalie, and Esther groups, reflecting the higher frequency of interactions within this group compared to the Marco and Julian cluster. The visualization also captures subtle relationship dynamics, such as Esther’s unidirectional engagement with William, which influences the latent space positioning of both Isabel and William through second-order interaction effects. These examples demonstrate the advantages of latent vector analysis in visualizing complex network data. By projecting interactions into a low-dimensional space, this approach uncovers hidden structures, including higher-order interactions and bridging relationships, while precisely representing both directionality and intensity of student connections. This methodology provides valuable insights into group dynamics, particularly for applications in educational research.
3.5.2 Network Mediation Model Results
Our investigation extends to examining how interaction patterns influence students’ mathematical achievement through network mediation analysis. This analytical approach specifically evaluates how network structures mediate the relationship between gender and mathematics learning outcomes. The network mediation analysis quantifies the three key components: the natural indirect effect (NIE), the natural direct effect (NDE), and the total effect (TE). Tables 8 and 9 present the comprehensive results of these analyses.
Post.M Post.SD 2.5% 97.5%
NIE 0.7758 1.2468 -1.7182 3.3649
NDE -5.475 1.5948 -8.8699 -2.48 *
TE -4.6992 0.8321 -6.4324 -3.056 *
Table 8: Network Mediation Model Results for EOI Network
Analysis of the EOI network, presented in Table 8, indicates that NDE is statistically significant and negative, while NIE through the network is not statistically significant. This negative NDE after controlling for proficiency of pre-test as a confounder, suggests that gender has a clear negative direct effect on post-test scores, meaning female students tend to achieve lower math scores independent of the mediating EOI network structure.
Post.M Post.SD 2.5% 97.5%
NIE 2.8653 1.3964 0.5089 5.9881 *
NDE -6.9407 1.9286 -11.0408 -3.3803 *
TE -4.0754 1.1155 -6.344 -1.7559 *
Table 9: Network Mediation Model Results for EXP Network
In contrast, for the EXP network, Table 9 reveals a different pattern. The negative NDE, which is consistent with the EOI network, shows that female students’ scores tend to be lower when considering gender independently of network effects. However, the positive NIE suggests that gender influences network structure in ways that enhance female students’ performance. These findings highlight the crucial role of network interactions, particularly within the EXP network, in mediating and potentially mitigating gender-based performance differences through collaborative learning processes.
4 Discussion
In the current study, we proposed the analysis of classroom interactions from their dialogues through an innovative combination of prompt engineering and network analysis methodologies. The prompt engineering approach enables automated classification of interaction patterns without requiring extensive pre-labeled datasets, significantly improving upon traditional manual labeling methods in both efficiency and scalability. Building on this automated classification, our network analysis framework represents students as nodes and their interactions as edges, providing a comprehensive visualization and analysis of classroom information flow and engagement patterns. Through the projection of these networks onto latent spaces, we quantified structural characteristics that provide the understanding structure of classroom interactions.
The network mediation analysis provides important insights into the relationship between classroom dynamics and student achievement. Our findings demonstrate that network mediation models possess unique capabilities for capturing intricate social patterns in educational environments which traditional analytical approaches often fail to identify due to the multidimensional nature of classroom interactions. This research opens several promising directions for future research, particularly in the incorporation of additional variables and the development of more sophisticated network representations, such as dynamic networks to capture temporal evolution of interactions or implement multilayer networks to represent diverse forms of student engagement simultaneously. Our methodological integration validates the effectiveness of combining prompt engineering with network analysis for understanding how student interactions influence learning outcomes.
Our findings suggest several promising directions for future research to deepen our understanding of classroom dynamics and refine analytical methodologies. A key area for expansion involves the analysis of teacher-student interactions, which we did not address in the current study. Teachers serve as essential facilitators of learning, significantly influencing classroom dynamics and student engagement patterns. Examining these teacher-student interactions would provide valuable insights into how instructional practices shape educational outcomes, particularly in the development of critical thinking skills and collaborative learning. Such analysis could inform the optimization of teaching strategies and contribute to enhanced classroom effectiveness.
Future research could enhance the accuracy of interaction analysis by exploring model fine-tuning with high-quality labeled datasets, extending beyond our current prompt engineering approach. Such fine-tuning could significantly improve the model’s precision and generalizability in analyzing complex educational interactions. Additionally, the systematic integration of expert knowledge and carefully curated datasets into the prompt engineering process would strengthen the methodological reliability and validity. This enhanced approach could provide more nuanced insights into classroom interaction patterns while maintaining the efficiency advantages of automated analysis.
A critical direction for future research involves developing automated prompt-generation methodologies. The current reliance on human judgment for prompt creation and selection introduces potential variability and bias into the analytical process. An automated approach that selects prompts based on their likelihood of generating desired outputs would enhance both the efficiency and consistency of the labeling process. Such automation would transform this methodology into a more scalable and reliable tool for educational research, enabling broader application across diverse educational contexts while maintaining analytical rigor.
Acknowledgement
We thank the editor, associate editor, and reviewers for their constructive comments. This work was partially supported by the National Research Foundation of Korea [grant number RS-2023-00217705 and RS-2024-00333701; Basic Science Research Program awarded to IHJ] and the ICAN (ICT Challenge and Advanced Network of HRD) support program [grant number RS-2023-00259934], supervised by the IITP (Institute of Information & Communications Technology Planning & Evaluation). Correspondence should be addressed to Ick Hoon Jin, Department of Applied Statistics, Department of Statistics and Data Science, Yonsei University, Seoul, Republic of Korea. E-mail: ijin@yonsei.ac.kr.